OpenStudy (anonymous):

There are 6 songs on a CD. The CD player is programmed to play 6 songs in random order without repeats. What s the probability it will play songs in order listed on the CD? A) 1/6 B) 1/36 C) 1/720 D) 1/46,656

OpenStudy (anonymous):

a

OpenStudy (anonymous):

can i have my medal now -_-

OpenStudy (anonymous):

how did you get it? :O

OpenStudy (anonymous):

its too easy to explain lol

OpenStudy (anonymous):

lol -1 for that

OpenStudy (anonymous):

The answer should be (c) i guess

OpenStudy (anonymous):

its not A bro

OpenStudy (anonymous):

its C => 1/6! = 1/720

Directrix (directrix):

1/6 * 1/5 * 1/4* 1/3 *1/2 *1 = 1/ 6! = 1/720

OpenStudy (anonymous):

-1 for marsbar

OpenStudy (anonymous):

sorry misread the question myfault thought it wanted to no the probablity it will play all songs lol im a dumby thats why i dont answer questions lol

OpenStudy (anonymous):

so it's not C? O_o

OpenStudy (anonymous):

idk ask the geniuses cuz im out peace!!!!

OpenStudy (anonymous):

its C, there are 6! number of ways to play it, so the probability of playing it 1 particular way would be 1/6!

OpenStudy (anonymous):

1/6 or 1/720 im so confused now =P

OpenStudy (anonymous):

so its a?

OpenStudy (anonymous):

1/6! =>6! = 6 factorial -> ! = factorial

OpenStudy (anonymous):

6 factorial = 6! = 720

OpenStudy (anonymous):

you got me lost and im not even the one that asked the question lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

LOL thanks everyone! i got it.

Directrix (directrix):

C) 1/720 --> ANswer 1/6 * 1/5 * 1/4* 1/3 *1/2 *1 = 1/ 6! = 1/720